please help me solve this with steps on how to do these problems.. I don't really understand where to begin..
the instructions state:
Find the EQUATION of the line with the given slope and passing through the given point. Your answer should be in slope-intercept form: y = mx + b
the problem:
m = 2/3; (3,-3)

Answer by ad_alta(240)

Since we know the slope is 2/3, all we need is 'b.' Therefore y=(2/3)x+b. We know (3,-3) is on the line, so it must satisfy the equation. Thus, -3=(2/3)(3)+b gives b=-5. The equation is y=(2/3)x-5. That's all there is to it!